# Quantization Effects on Neural Networks Perception: How would quantization change the perceptual field of vision models?

### v1

@@ 3.2 Qualitative Results
Figure 2 depicts an example of an image from the study dataset with the related overlaid saliency map. It also depicts the different CAMs associated with different models and levels of precision.
$VGG-16$, $ResNet50$, and $DenseNet121$ exhibit a low level of impact in terms of changes on CAMs, at the int16 precision. On the other hand, we notice significant shifts when comparing at the int8 quantization level. Which is consistent with the deductions from the empirical results.
Smaller networks like $MobileNet$, and $SqueezeNet$ show fewer changes across lower levels. We can especially notice that CAMs from $SqueezeNet$ remain unchanged.
$EfficientNet$ displays the highest level of changes across the models which means that it is much less robust. This is due to the simplicity of the network architecture and the small number of parameters.
@@ 4 Conclusion
This study investigated how quantization affects different neural network architectures using Class Activation Maps (CAMs) to gauge their alignment with human perception. We found that quantization with int16 precision strikes a good balance between efficiency and maintaining human-like perception (measured by SIM, CC, KLD). Notably, MobileNet and SqueezeNet exhibited exceptional robustness due to their efficient designs. Conversely, EfficientNet suffered more under quantization.
These findings offer valuable insights for deploying neural networks on resource-constrained devices. By understanding how different architectures respond to quantization, researchers and practitioners can make informed decisions to achieve a balance between model performance, interpretability, and efficiency in real-world applications.

### v2

@@ 3.2 Qualitative Results
Figure 2 depicts the impact of quantization on CAMs for the chosen deep learning models. Each row represents a different model (DenseNet-121, EfficientNet-B0, MobileNetV2, ResNet-50, SqueezeNet-1.0, VGG-16), and each column represents the effect of a different quantization level (f32, int16, int8). The first column shows the original image and the second a ground truth segmentation of the salient object.
VGG-16 demonstrates a gradual shift in the perceptual field, where with the decrease of the quantization level, it expands its perceptual field, thus introducing more information noise about other parts of the image.
ResNEt-50 shows more robustness at the int 16 quantization level, but a significant change in the perceptual field areas at the int8 level.
DenseNet-121 shows similar behavior as ResNEt-50, but we notice that it tends to use a larger perceptual field covering more area of the stimuli.
MobileNet-v2 shows less robustness to quantization effects, where the perceptual field changes at all precision levels. We also notice that the network tends to have a larger and more uniform perceptual field. It also tends to heavily emphasise on larger areas of the image as shown by the heatmap.
In accordance with the empirical data, SqueezeNet-1.0 is the least affected by the quantization, we can notice that it tends to have uniforme and more moderate coverage of the image maximizing the amount of innformation gathered, this explains it’s hardiness against the loss of information stemming from quantization.
The EfficientNet-B0 network shows the least robustness in accordance with the empirical data. The shift in the perceptual fields across different quantization levels is evident through the figure.
In summary, our comprehensive analysis reveals the profound impact of quantization on the perceptual fields of various neural network architectures, where both quantitative and qualitative analysis align their findings.
@@ 4 Conclusion
This study investigated how quantization affects different neural network architectures using Class Activation Maps (CAMs) to gauge their alignment with human perception and the semantic content of the stimulus.
The study demonstrates that while some models, such as SqueezeNet-1.0, exhibit remarkable robustness to quantization, others, like EfficientNet-B0, show significant sensitivity, particularly at lower precision levels. These findings underscore the importance of considering quantization effects when deploying deep learning models on resource-constrained devices especially in use cases of interpretability.
These findings offer valuable insights for deploying neural networks on resource-constrained devices. By understanding how different architectures respond to quantization, researchers and practitioners can make informed decisions to achieve a balance between model performance, interpretability, and efficiency in real-world applications. Nonetheless, we need to mention that these results are still limited and that it would be beneficial to expand these results to more diverse datasets, models, domains, and tasks.